OpenStudy (anonymous):

quotient function (x^3 + x / x^2 - 1) and I need to find the derivative and stationery points. here is the derivative: (3x^2+1)(x^2-1) - 2x^2(x^2+1) (top line) / (x^2-1)^2 Now I need to solve for x but I am stuck.... please help!

OpenStudy (anonymous):

Do you mean (x^3 + x)/(x^2 - 1) or (x^3 + x/x^2 -1)?

OpenStudy (ranga):

top line last after the minus sign should be: 2x(x^3 + x)

OpenStudy (ranga):

\[f(x) = \frac{ x^3+x }{ x^2-1 }\]\[f'(x) = \frac{ (3x^2+1)(x^2-1) - 2x(x^3+x) }{ (x^2-1)^2 } = 0\]